OpenStudy (anonymous):

In a word game, you choose a tile from a bag, replace it, and then choose another. If there are 21 vowels and 15 consonants, what is the probability you will choose a consonant and then a vowel? 35/8 1/36 35/4 35/144

OpenStudy (anonymous):

If you replace the tiles, it should be the probability of drawing the consonant multiplied by the probability of drawing a vowel. (15/36)*(21/36) = (5/12)*(7/12) = (35/144) which is about 24.3055556% chance

OpenStudy (anonymous):

If you don't replace the tiles then instead do (15/36)*(21/35) do get 1/4 or 25%